OpenStudy (anonymous):

another... sqrt(z+1)+sqrt(z)=2

OpenStudy (anonymous):

\[\sqrt{z+1}+\sqrt{z}=2\]

OpenStudy (anonymous):

I got z=9/16 correct?

OpenStudy (anonymous):

Yeah 9/16 is correct...

OpenStudy (anonymous):

awesome! thank you! :)

OpenStudy (kainui):

Plug it in, do you get 2? Yep, so you're golden.